### v2

# Basic syntax from speech: Spontaneous concatenation in unsupervised deep neural networks

Abstract
Computational models of syntax are predominantly text-based. Here we propose that the most basic syntactic operations can be modeled directly from raw speech in a fully unsupervised way. We focus on one of the most ubiquitous and elementary properties of syntax—concatenation. We introduce spontaneous concatenation: a phenomenon where convolutional neural networks (CNNs) trained on acoustic recordings of individual words start generating outputs with two or even three words concatenated without ever accessing data with multiple words in the input. We replicate this finding in several independently trained models with different hyperparameters and training data. Additionally, networks trained on two words learn to embed words into novel unobserved word combinations. To our knowledge, this is a previously unreported property of CNNs trained in the ciwGAN/fiwGAN setting on raw speech and has implications both for our understanding of how these architectures learn as well as for modeling syntax and its evolution from raw acoustic inputs.
keywords:
language evolution , language acquisiton , concatenation , syntax from phonetics , unsupervised deep learning , generative adversarial networks , deep language learning
journal: Journal
[…]
[label1]organization=Department of Linguistics, University of California, Berkeley,country=United States
[…]
[label2]organization=Department of Computer Science, Iowa State University,country=United States
1 Introduction
How language evolved and what is required for humans to acquire syntax are among the central questions in linguistics and cognitive science (Christiansen and Kirby, 2003; Nowak and Komarova, 2001; Fitch, 2000). Due to lack of direct evidence for language evolution, many studies employ evolutionary modeling (Christiansen and Dale, 2003; O’Grady and Smith, 2018) in order to gain insights into the putatively most difficult problem in science. While many studies achieved substantial progress (Nowak and Krakauer, 1999; Kirby, 2000; Griffiths and Kalish, 2007; Zuidema and de Boer, 2009; Perfors and Navarro, 2014; O’Grady and Smith, 2018), most models, however, are either theoretical or operate with highly simplified and abstract assumptions. In this paper, we model the first step in the evolution/acquisition of syntax—concatenation—with unsupervised deep neural networks that are among the most realistic models of human language learning and learn linguistic representations from raw spoken language inputs (Beguš, 2021a).
[…]
As our main model of learning we adopt the ciwGAN/fiwGAN model (Beguš, 2021a) which model language as informative imitation and features several advantages over competing computational model and represents one of the closest approximations of human language learning using deep neural networks. The sounds of human speech are a measurable, physical property, but they also encode abstract linguistic information such as syntactic, phonological, morphological, and semantic properties. It has been shown that the model learns symbolic-like rule-like linguistic representations from raw speech (Beguš, 2021a, b, 2020) which illustrates that discrete mental representational units can emerge from continuous physical properties in deep neural networks. The model features several features of human spoken language learning that other models lack: learning by imitation, learning through the production-perception loop, and communicative intent.
Prior models of human syntax are predominantly text-based and mostly lack explicit mechanisms that can be directly paralleled to language processing in the human brain. Hearing human learners, however, acquire syntax from acoustic inputs and process it with biological neural mechanisms. Few studies model syntax from speech with artificial neural networks. Previous work has shown that speech-trained models can feature syntactic representations, but the models featured specific syntactic mechanisms (Lai et al., 2023) or were pre-trained (Singla et al., 2022). These architectures do not model the emergence of syntax.
Here, we propose that we can model a shift from single unit system to a concatenative system using ciwGAN. We model how compound signals or concatenated words can arise spontaneously in deep neural networks trained on raw speech in a fully unsupervised manner. We thus also provide evidence suggesting that domain-general learning models (such as CNNs) spontaneously concatenate signals in a learning setting that involves learning by imitation and production-perception learning from raw spoken language data (ciwGAN). The paper thus introduces a model of the first step in the development from holistic non-syntactic signals to concatenated signals that represents the most elementary syntactic processing. We also show that the observed spontaneous concatenation is not an idiosyncratic property of a single model, but emerges in several replicated models as well as in models that features several changes in the architecture.
2 Methods
To test whether CNNs can spontaneously concatenate lexical items, we conduct two sets of experiments. In the first set of experiments (three one-word experiments), we train the networks on single-word inputs and test the networks for concatenated outputs. In the second experiment, we train the networks on one-word and two-word inputs (the two-word experiment) and withhold a subset of two-word combinations. We then test whether words can be embedded into novel unobserved combinations in the output. Such a design also mimics one-word and two-word stages in language acquisition (Berk and Lillo-Martin, 2012).
Because the networks are trained in the GAN setting, the Generator never accesses the training data directly, but generates innovative outputs. It has been shown that GANs innovate in highly interpretable ways that produce novel words or sound sequences (Beguš, 2021a). Here we test whether innovations can produce spontaneously concatenated words.
2.1 The model
We train the ciwGAN and modified fiwGAN models (Beguš, 2021a). ciwGAN/fiwGAN models are information-theoretic extensions of GANs (based on InfoGAN, WaveGAN and DCGAN; Chen et al. 2016; Donahue et al. 2019; Radford et al. 2015) designed to learn from audio inputs.
The ciwGAN/fiwGAN architectures involve three networks (Figure 1 and Table 1): the Generator that takes latent codes $c$ (either one-hot vectors or binary codes) and random latent space variable $z$ ($z\sim\mathcal{U}(-1,1)$) and through five or six upconvolutional layers generates 1.024s or 2.048s audio with 16 kHz sampling rate (16,384 or 32,768 samples). The audio is then fed to the Discriminator, which evaluates realness of the output via the Wasserstein loss (Arjovsky et al., 2017). The unique aspect of the ciwGAN/fiwGAN architecture is a separate Q-network, which is trained to estimate the Generator’s hidden code $c$. In the ciwGAN architecture, the latent code $c$ is a one-hot vector; in the fiwGAN architecture the latent code is a binary code. It has been shown that even if there is a mismatch between the number of lexical items and the number of classes, lexical learning nevertheless occurs in the models (Beguš, 2021a).
Figure 1: The architecture of ciwGAN used in the two-second one-word experiment.
Generator Discriminator Q-Network
Layer Dimension Layer Dimension Layer Dimension
$z$ 100 $\times$ 1 input 32768 $\times$ 1 input 32768 $\times$ 1
fc + reshape 16 $\times$ 2048 conv0 8192 $\times$ 64 conv0 8192 $\times$ 64
upconv0 64 $\times$ 1024 conv1 2048 $\times$ 128 qconv1 2048 $\times$ 128
upconv1 256 $\times$ 512 conv2 512 $\times$ 256 qconv2 512 $\times$ 256
upconv2 1024 $\times$ 256 conv3 128 $\times$ 512 qconv3 128 $\times$ 512
upconv3 4096 $\times$ 128 conv4 32 $\times$ 1024 qconv4 32 $\times$ 1024
upconv4 16384 $\times$ 64 conv5 16 $\times$ 2048 qconv5 16 $\times$ 2048
upconv5 32768 $\times$ 1 flatten + logit 1 $\times$ 1 flatten + logit 1 $\times$ 1
Table 1: The structure of the Generator, the Discriminator, and the Q-network in the two-second experiments (based on Donahue et al. 2019; Beguš 2021a).
During training the Generator learns to generate data such that it increases the Discriminator’s error rate and decreases the Q-network’s error rate. In other words, the Generator needs to learn to encode unique information into its acoustic inputs, such that the Q-network is able to decode unique information from its generated sounds. The training between the Generator and the Q-network mimics the production-perception loop in speech communication: after training, the Generator learns to generate individual words given a latent code $c$ (Beguš, 2021a) and the Q-network learns to classify unobserved words with the same corresponding codes (Beguš and Zhou, 2022). Since learning is completely unsupervised, the Generator could in principle encode any information about speech into its latent space, but the requirement to be maximally informative causes it to encode linguistically meaningful properties (both lexical and sublexical information; Beguš and Zhou 2022). Such a setting not only replicates the production-perception loop, but is also one of the few architectures featuring traces of communicative intent (between the Generator and the Q-network). Unlike in generative models trained on next sequence prediction or data replication where no communicative intent exists, the training objective between the Generator and the Q-network is to increase mutual information between the latent space and the data such that the Q-network can retrieve the information (latent code) encoded into the speech signal by the Generator.
CiwGAN and fiwGAN have been shown to be highly innovative in linguistically interpretable ways (Beguš, 2021a, b). For example, the Generator produces new words or new sound sequences that it never accesses during training. Crucially, the Generator never directly accesses the data: it learns by generating data from noise such that the Discriminator fails to distinguish real and generated data. In this respect, it mimics learning by imitation in human language (rather than replication as is the case with variational autoencoders).
Three models are trained for the purpose of this paper, and one model is pretrained from Beguš (2021a). The four models were trained for 16 hours on a single GPU (NVIDIA 1080ti). For guidelines on the choice of the number of steps during training, see Beguš (2020). We use the TIMIT (Garofolo et al., 1993) database for training. The number of parameters is given in A. We take the standard hyperparameters (from Donahue et al. 2019 and Beguš 2021a). Because the outputs are salient and rarely ambiguous, we combine transcriptions performed by the authors with transcriptions performed by Whisper (Radford et al., 2022). Generated audio files and models’ checkpoints are available at the anonymous link: https://osf.io/przuq/?view_only=9d19a26f0bb84a3ea4db8e6844b37985.
2.2 Data
The training dataset consists of sliced lexical items from the TIMIT database of spoken English (Garofolo et al., 1993), such that each item is a single spoken word.
In order to test robustness of spontaneous concatenation across different models and settings, we conduct four experiments with varying parameter choices. We alter the number of training steps, training words, word-types, and tokens, randomness of padding, length of output and model type. The experiments are summarized in Table 2.
In Experiment 1, the one-second one-word experiment, we test a pre-trained model from Beguš (2021a) for signs of concatenation. As specified in Beguš (2021a), the 5-layer Generator outputs only 1.024s of audio and data is never left-padded (only right padded) which controls for the effect of padding on concatenation. The model is trained using 5 lexical items with  600 tokens each: oily, rag, suit, year and water. The smaller number of lexical items chosen for training is advantageous from the perspective of modeling language acquisition as it mimics the stage when language-acquiring children have fewer lexical items in their inventory.
In Experiment 2 (simple two-second one-word), we replicate the results with another one-word experiment trained on box, greasy, suit, under, and water. Here, each item is represented by a single token (to test how the model performs with less variability during training) and randomly padded with silence to a length of 2s to produce 100 distinct data points for each class, for a total of 500 data points used in training (the two-second one-word experiment).
We perform Experiment 3 (complex two-second one-word) by training a model where each lexical item is represented by approximately 600 distinct tokens, as found in the TIMIT dataset, to test the ability of the model to learn in a more complex environment. Again, for this experiment each data sample is randomly padded with silence to a length of 2s.
Experiment Words per file Length of output Padding Types Tokens Steps Model
1 1 1.024s right 5 $\sim$600 8,011 ciwGAN
2 1 2.048s random 5 1 8,956 ciwGAN
3 1 2.048s random 5 $\sim$600 18,842 ciwGAN
4 2 2.048s random 3 1 9,166 fiwGAN
Table 2: An overview of parameters used in the four experiments.
In Experiment 4 (two-second two-word), we use 3 lexical items with a single token each: greasy, suit, and water. 100 data points each 2s in length are generated in an analogous process to the first experiment, but for each combination of two items (i.e. greasy, suit, and water alone, greasy followed by water, water followed by greasy etc.). However, we withhold the combination suit/greasy, such that they do not appear together in the training set in any order, to produce a final training set of 700 data points.
For the one-word experiments, we use the ciwGAN model with one-hot encoding and five levels, such that each of the five unique words can be represented with a unique one-hot vector. In the two-word experiment, we use a modified fiwGAN (binary code). The binary code is limited to three bits, but each code can have up to two values of 1 (e.g. [1,0,0] and [1,1,0]).
3 Results
To test whether the models can spontaneously concatenate, the networks are trained for 8,011 (pretrained one-second one-word), 8,956 (simple two-second one-word), 18,842 (complex two-second one-word), and 9,166 (two-word fiwGAN) steps and analyze generated data. We use the technique proposed in Beguš (2020) to analyze the relationship between linguistically meaningful units and latent space. According to this technique, setting individual latent space variables to values outside of the training range reveals the underlying linguistic value of each variable.
3.1 One-word model
In the one-second one-word model (Experiment 1 in Table 2), the Generator learns to associate each unique one-hot code with a unique lexical item in a fully unsupervised and unlabeled manner (Beguš, 2021a). The Generator’s input during training is a one-hot vector with values 0 or 1. For example, the network learns to represent suit with $[1,0,0,0,0]$. To test this observation, we set the one-hot vector to values outside the training range (e.g. $[5,0,0,0,0]$), which overrides lower-level interactions in the latent space (Beguš, 2020). This causes the Generator to output suit at near categorical levels (9 times out of 10), revealing the underlying value of the code. This further reveals that $[0,1,0,0,0]$ encodes year (8 times out of 10) and $[0,0,1,0,0]$ encodes water (10 times out of 10).
In addition to lexical learning, we observe a robust pattern in the pretrained one-word one-second model: the networks trained on one-word inputs generate two-word outputs when the one-hot values are set to negative values outside of the training range. For example, when the latent code is set to $[0,-2,-2,-2,0]$, the Generator consistently outputs a two-word output suit year (8 times out of 10). For $[-3,-2,-3,-2,2]$, the network consistently outputs rag year (8 times out of 10; Figure 2). These concatenations occur despite the fact that the training data is always left-aligned, the Generator never accesses the data directly and the Discriminator only sees single words.
Figure 2: The suit year (left) output and the rag year (right) from the one-second one-word model. All spectrograms are created in Praat (Boersma and Weenink, 2015).
To show that this outcome is not an idiosyncratic property of one model and that it is indeed the negative values that encode concatenated outputs, we also analyze the simple two-second one-word model (Experiment 2 in Table 2), which was trained for the purpose of this experiment. The inputs to the Discriminator in this case are also single words only, but they are longer (2s) and randomly padded with silence on the left and right. While high positive values occasionally yield two-word outputs in this model, negative values are consistently associated with two-word outputs. For example, $[-50,-50,0,-50,0]$ (with extreme values) consistently encodes box greasy (9 times out of 10), and $[-50,-50,-50,0,0]$ consistently encodes greasy under (10 times out of 10). Positive values of the same codes with extreme values (50) produce completely unintelligible outputs (noise).
The same outcomes are observed in the Experiment 3 (in Table 2) where the models are trained on approximately 600 token for each lexical item (for a quantitative analysis of this experiment, see Section 3.2 and Figure 4).
In addition to several two-word concatenated outputs, the network even occasionally generates a three-word concatenated output box under water (Figure 3) for the latent code $c$ with all negative values $[-3,-1,-1,-1,-1]$ (2 times out of 10).
Figure 3: The three-word concatentated output box under water. Independently, the second word (under) is somewhat difficult to analyze, but given only five training words, it is clearly the closest output to under.
3.2 Statisical analysis
Figure 4: The average number of words generated by the two one-word two-second models are plotted as a function of the first 3 bits of the latent code. The remaining two bits of the latent code are maintained at a value of 0. As in the other trials, each bitstring was tested with 10 sets of latent space values.
Raw data in Figure 4 shows that the simple one-word two-second model generates between 0 and 3 words for each z vector, generating substantially more words at negative bit values. To test the observation that negative unobserved values correspond to concatenated outputs, we systematically generate samples where each of the 5 bits is interpolated from the set of values $\{-6,-3,0,3,6\},$ and all permutations of bits are tested ($5^{5}$). Each permutation is tested with 10 sets of latent space values (constant across different permutations), for a total of $31,250$ samples. We automatically annotate our outputs using a pretrained Whisper-small model (Radford et al., 2022), which we fine-tune using a dataset consisting of 400 model outputs manually annotated by the authors. Whisper is used to annotate for single (failure; 0) vs. multiple (success; 1) word outputs. We estimate Whisper’s accuracy on labeling single- and multiple-word outputs at 88% (based on author’s annotations; Table 4). The data were fit to a logistic regression mixed effects model (details of the model in Figure 5). There is a clear and significant negative relationship between the sum value of bits and the probability of a multiple-word output ($\beta=-0.09,z=-54.42,p<0.0001$; Figure 5). The model with interactions between individual bits also shows that negative values correspond to significantly higher proportion of two-word outputs (Figure 5, Table 3).
We perform a similar analysis on the complex one-word two-second model (Experiment 3 in Table 2), where we systematically generate samples where each of the 5 bits is interpolated from the set of values $\{-4,-2,-1,0,1,2,4\},$ and all permutations of bits are tested $(7^{5}),$ Again, each permutation is tested with 10 sets of latent space values (constant across different permutations), for a total of $168,070$ samples. All samples were subsequently transcribed using the same fine-tuned Whisper model used in the previous experiments, and the results suggest a similar correlation between negative code values and the proportion of two-word outputs (Figure 4).
The data of this experiment were fit to a logistic regression model which again shows a significant relationship between the proportion of multiple-word outputs (the dependent variable) and the sum of all the latent code bits (the independent variable) ($\beta=-0.14,z=-167.5,p<0.0001$). All details of the model and plotted estimates are in Figure 8.
Figure 5: (top) Predicted values of the logistic regression mixed effects model with the proportion of two-word outputs (one-word output = failure, two-word output or more = a success) as the dependent variable and sum of bits 1–5 as the predictor with 95% confidence limits (Experiment 2 in Table 2). Outputs with no transcribed words were removed. The random effect structure involves random intercepts for each of the 10 unique random latent space samples $z$ as well as random slope for sum of bits. The values were obtained with the effects package (Fox and Weisberg, 2019; Fox, 2003). (bottom) Predicted values for one (bit1 * bit2 * bit5) of the many interactions of the logistic regression mixed effects model with the proportion of two-word outputs (one-word output = failure, two-word output = a success) as the dependent variable and individual bits with all interactions (including the five-way interaction) as the predictors with 95% confidence limits. The random effect structure involves only the random intercept for each of the 10 unique random latent space samples $z$. The values were obtained with the effects package (Fox and Weisberg, 2019; Fox, 2003). Model estimates are given in Table 3. Not all interactions show the same effect.
3.3 Two-word model
In the two-word experiment, the models get one-word and two-word inputs (thus mimicking the two-word stage in language acquisition). The models are only trained on three words and their combinations, except for the suit/greasy combination withheld.
In the fiwGAN two-word model (Experiment 4 in Table 2), the Generator consistently outputs the unobserved greasy suit for $[15,0,0]$ (17 times out of 20), which suggests the network learned this unobserved combination as one of the possible sequences and encoded it with a one-hot value. To evaluate the presence of the withheld greasy suit pair systematically, we conduct an exploration of the latent space (all permutations where each variable takes on values in $\{-13,-8,-3,-2,-1,0,1,2,3,8,13\}$) in Figure 6. The withheld suit/greasy combinations occur most often for codes around [-10,0,0], or otherwise when at least one bit is negative (Figure 6), suggesting that the model has associated the unobserved suit/greasy combination with that code. It appears that the negative values of the latent code $c$ again encode unobserved novel combinations.
Figure 6: The frequency of observed suit/greasy pairs is plotted as a function of the latent code, for the fiwGAN two-word model. Samples were generated by varying each bit between the values [-13, -8, -3, -2, -1, 0, 1, 2, 3, 8, 13], taking all possible permutations ($11^{3}$). Each bitstring was tested with 10 sets of latent space values, and annotations were performed automatically using Whisper.
3.4 Repetition
In addition to two-word concatenation and embedding of words into novel combinations, we also observe outputs with repeated words in all our trained models. The training data never includes the same word repeated, yet the models frequently include repeated words. For example, the two-second one-word model consistently outputs greasy greasy for $[0,0,-40,0,0]$ (7 times out of 10; Fig. 7). Of the 31,250 samples annotated in the analysis, 2753 (8.8%) contained repeated words, suggesting that the model systematically learned this behavior. This is significant because repetition or reduplication is one of the most common processes in human language and language acquisition (Berent et al., 2016; Dolatian and Heinz, 2020). Additionally, full or total reduplication (where the entire word is repeated) is among the most computationally complex morphophonological processes (Dolatian and Heinz, 2020) because it represents unbound copying at the segmental level. It has been shown elsewhere that deep convolutional neural networks can learn partial reduplication (where only a part of the word is repeated) from raw speech and extend it to novel unobserved data (Beguš, 2021b). Our results suggest that total reduplication (or unbound copying; Dolatian and Heinz 2020) can arise spontaneously in these models.
Figure 7: The suit greasy water output from the two-second two-word model (top). The greasy greasy output from the two-second one-word model (bottom).
3.5 Single batch model
We consider the possibility that spontaneous concatenation arises solely due to the models being trained in a batched training configuration, where different samples in the same batch could be concatenated when their gradients are combined into a single training update. To demonstrate that spontaneous concatenation occurs even in the absence of batching, we train an additional one-word model using the same dataset and model parameters as the complex two-second one-word model, but with the batch size set to $1$ We trained this model for 41,922 steps. It is well-known that batch size can affect training outcomes (Lin, 2022). Because of the small batch size, the words generated by this model are qualitatively poorer which prevents an automated Whisper analysis of the outputs. However, we continue to consistently observe multiple-word outputs in the negative latent space, indicating that this behavior is not solely attributable to batching. We randomly generated 10 examples from this model with the latent code set at [-1, -1, -1, -1, -1]. All 10 examples feature two or more word sequences, although the audio quality is not as high in the higher batch models. Waveforms and spectrograms of the ten examples are given in Figure 9.
4 Discussion
4.1 Limitations and future directions
This paper models concatenation of acoustic lexical items. Syntax is substantially more complex than concatenation (Berwick and Chomsky, 2019). Exploration of other syntactic properties as well as of word-level compositionality in these models is left for future work. That compositionality on the sublexical level occurs in these models has been shown elsewhere (Beguš, 2021a; Beguš and Zhou, 2022). Our work focuses on modeling the stage of language evolution/acquisition where the signals become compounded but not yet fully compositional. This stage has been described in Luuk and Luuk (2014). We also train the network on a relatively small number of lexical items (5) and a small number of tokens (from one to about 600). The small number of lexical items is representative of the earliest stages of language acquisition when the number of lexical items is highly limited (Bates et al., 1994).
4.2 Why negative values?
The Generator makes use of the unutilized space in the latent code to encode unobserved but linguistically interpretable outputs. During training, the Generator is trained on only two values of the latent code: 0 and 1. In the one-second one-word model, individual codes represent unique individual words, which suggest lexical learning emerges in the positive values in these models (Beguš, 2021a). The network never accesses two-word inputs and it never gets negative values in the latent code $c$ during the training. It appears that the network is biased to concatenate and that it uses the unobserved latent code space to encode unobserved concatenated outputs.
4.3 Why concatenation?
Our results are at first glance surprising: there are seemingly no reasons for the Generator to produce concatenated words. The training data never provides evidence for concatenated outputs. It should be easy for the Discriminator to distinguish single- from multiple-word outputs and estimate multiple-word outputs as diverging from the real data distributions. One possible reason for the spontaneous concatenation in GANs is that unlike any other architecture, GANs learn primarily by imitation and they do not replicate data, but rather learn to construct data by innovation. In other words, in the GAN architecture, the agent that generates data (the Generator) never accesses training data directly, but learns to construct data from noise in the latent space by generating data such that another agent is unable to distinguish it from training data - effectively forcing the imitative principle. Imitation is a prominent facilitator of language acquisition in humans, especial in early spoken language learning (Clark, 1977; Kuhl and Meltzoff, 1996; Masur and Rodemaker, 1999; Rasilo and Räsänen, 2017).
It is possible that the imitative and innovative design of GANs mimics the necessary conditions for the switch from simple to compound signals in humans. Learning by imitation and innovation is a crucial condition in human language. It is possible that based on these same conditions, human language learners start uttering simple signals as concatenated and analyzing them as complex signals. It is also possible that similarly to encoding novel unconcatenated outputs in the unobserved part of the latent space, the brain encodes newly concatenated phrases as the brain expands during language acquisition. The negative latent space unobserved during training can, from a very abstract, modeling perspective, be paralleled to growing neuronal connections during speech acquisition. That syntactic processing develops at a relatively late stage and requires developments in the brain has been shown before (Hahne et al., 2004; Skeide et al., 2014). It has been shown elsewhere that the internal representations in ciwGAN/fiwGAN models closely match brain responses to sound of language in raw untransformed form (Beguš et al., 2023). These assumptions are, however, highly speculative at this point and require further evaluation.
5 Conclusion
Modeling the evolution from holistic single-call units into a combinatorial syntax is not a trivial task. How human syntax emerged or how humans evolved from the stage of holistic calls into a fully concatenative and then compositional syntax is an unsolved problem. While other studies offer proposals about how this shift may have happened (Jackendoff, 1999; Luuk and Luuk, 2014; Progovac, 2015), our paper shows that the shift spontaneously occurs in computational models. Many studies exist that model evolution of syntax (Christiansen and Kirby, 2003), but the present study is, to our knowledge, the first that shows how concatenation emerges spontaneously in a fully connectionist model that did not have any explicit predispositions for concatenation and that is trained on raw data. We speculate that concatenation emerges because the models learn by imitation and innovation rather than replication of data (Section 4).
Our results suggest that the Generator network in the ciwGAN architecture not only learns to encode information that corresponds to lexical items in its audio outputs, but also spontaneously concatenates those lexical items into novel unobserved two-word or three-word sequences.
The ability of unsupervised deep neural networks trained on raw speech to concatenate words into novel unobserved combinations has far-reaching consequences. This means that we can model basic syntactic properties directly from raw acoustic inputs of spoken language, which opens up potential to model several other syntactic properties directly from speech with CNNs and GANs.
From the perspective of evolution of syntax, the results suggest that a deep neural network architecture with no language-specific properties can spontaneously begin generating concatenated signals from simple signals. The step from one-word stage to two-word stage is necessary both in evolution of human language as well as during language acquisition. Our second experiment mimics the two-word stage. We argue that unsupervised deep learning models not only concatenate single words into multi-word outputs, but are also able to embed words into novel unobserved combinations once the model is trained on multiple-word inputs.
Further research into the relationship between basic syntactic properties that spontaneously emerge in these fully unsupervised models trained on raw speech and the structure of the latent space has the potential to yield insights for the study of syntactic theory, language acquisition, and language evolution. By evaluating these models on syntactic properties of spoken language, we should also get a better understanding of computational limits of unsupervised CNNs trained on raw speech.
Finally, modeling syntax from raw speech with deep neural networks is informative not only for cognitive science and linguistics, but for machine learning research in general as well. Speech processing increasingly by-passes text (Lakhotia et al., 2021). Understanding syntactic capabilities of spoken language models can provide information for future architectural choices as natural language processing expands from text to raw audio modeling.
Acknowledgements
Author contributions
G.B.: Conceptualization, Data curation, Formal analysis, Funding acquisition, Investigation, Methodology, Project administration, Resources, Software, Supervision, Validation, Visualization, Writing – original draft, Writing – review and editing. T.L.: Conceptualization, Data curation, Formal analysis, Investigation, Software, Visualization, Writing – original draft, Writing – review and editing. Z.W.: Conceptualization, Data curation, Methodology, Writing – review and editing.
Conflict of interest
The authors declare no conflicts of interest.
Appendix A Appendix
Estimate Std. Error z value Pr($>$||||z||||)
(Intercept) -0.65 0.02 -36.7 0.0000
bit1 -0.21 0.00 -45.3 0.0000
bit2 -0.11 0.00 -24.4 0.0000
bit3 0.01 0.00 1.6 0.1013
bit4 -0.19 0.00 -41.2 0.0000
bit5 -0.17 0.00 -39.8 0.0000
bit1:bit2 0.03 0.00 30.5 0.0000
bit1:bit3 -0.02 0.00 -14.7 0.0000
bit2:bit3 -0.01 0.00 -7.0 0.0000
bit1:bit4 -0.00 0.00 -1.7 0.0911
bit2:bit4 0.03 0.00 28.2 0.0000
bit3:bit4 -0.02 0.00 -17.8 0.0000
bit1:bit5 0.02 0.00 22.7 0.0000
bit2:bit5 -0.03 0.00 -29.7 0.0000
bit3:bit5 -0.01 0.00 -7.2 0.0000
bit4:bit5 0.04 0.00 33.2 0.0000
bit1:bit2:bit3 -0.00 0.00 -13.9 0.0000
bit1:bit2:bit4 -0.01 0.00 -32.4 0.0000
bit1:bit3:bit4 0.01 0.00 39.1 0.0000
bit2:bit3:bit4 -0.00 0.00 -9.1 0.0000
bit1:bit2:bit5 -0.00 0.00 -1.9 0.0551
bit1:bit3:bit5 -0.00 0.00 -16.0 0.0000
bit2:bit3:bit5 0.00 0.00 0.5 0.5923
bit1:bit4:bit5 -0.00 0.00 -12.4 0.0000
bit2:bit4:bit5 -0.01 0.00 -19.8 0.0000
bit3:bit4:bit5 0.00 0.00 2.7 0.0060
bit1:bit2:bit3:bit4 0.00 0.00 2.5 0.0135
bit1:bit2:bit3:bit5 0.00 0.00 15.1 0.0000
bit1:bit2:bit4:bit5 0.00 0.00 17.8 0.0000
bit1:bit3:bit4:bit5 -0.00 0.00 -6.8 0.0000
bit2:bit3:bit4:bit5 0.00 0.00 3.5 0.0005
bit1:bit2:bit3:bit4:bit5 -0.00 0.00 -9.8 0.0000
Table 3: Estimates of the logistic regression mixed effects model with the proportion of two-word outputs (one-word output = failure, two-word output = a success) as the dependent variable and individual bits with all interactions (including the five-way interaction) as the predictors. The random effect structure involves only the random intercept for each of the 10 unique random latent space samples $z$.
Whisper annotated # of words
Human annotated # of words 1 2 3 Total
1 47 9 1 57
2 3 28 5 36
3 0 2 5 7
Total 50 39 11 100
Table 4: The accuracy of the Whisper speech-to-text model is assessed over 100 audio samples generated randomly using the two-second one-word ciwGAN model. Each sample was annotated manually by the authors and automatically using Whisper. For the classification of one word versus multiple words, Whisper agreed with the authors 88% of the time.
Figure 8: Predicted values of the logistic regression model with the proportion of two-word outputs (one-word output = failure, two-word output or more = a success) as the dependent variable and sum of bits 1–5 as the predictor with 95% confidence limits (Experiment 3 in Table 2). No random effects were included because mixed effects models result in singular fits. Outputs with no transcribed words were removed. The values were obtained with the effects package (Fox and Weisberg, 2019; Fox, 2003).
Figure 9: Ten randomly generated examples with the latent code [-1, -1, -1, -1, -1] of the single batch model.