### v3

# Basic syntax from speech: Spontaneous concatenation in unsupervised deep neural networks

Abstract
Computational models of syntax are predominantly text-based. Here we propose that the most basic first step in the evolution of syntax can be modeled directly from raw speech in a fully unsupervised way. We focus on one of the most ubiquitous and elementary suboperation of syntax—concatenation. We introduce spontaneous concatenation: a phenomenon where convolutional neural networks (CNNs) trained on acoustic recordings of individual words start generating outputs with two or even three words concatenated without ever accessing data with multiple words in the input. We replicate this finding in several independently trained models with different hyperparameters and training data. Additionally, networks trained on two words learn to embed words into novel unobserved word combinations. We also show that the concatenated outputs contain precursors to compositionality. To our knowledge, this is a previously unreported property of CNNs trained in the ciwGAN/fiwGAN setting on raw speech and has implications both for our understanding of how these architectures learn as well as for modeling syntax and its evolution in the brain from raw acoustic inputs. We also propose a potential neural mechanism called disinhibition that outlines a possible neural pathway towards concatenation and compositionality and suggests our modeling is useful for generating testable prediction for biological and artificial neural processing of speech.
keywords:
language evolution , language acquisition , concatenation , syntax from phonetics , unsupervised deep learning , generative adversarial networks , deep language learning
journal: tba
[…]
[label1]organization=Department of Linguistics, University of California, Berkeley,country=United States
[…]
[label2]organization=Department of Computer Science, Iowa State University,country=United States
1 Introduction
How language evolved and what is required for humans to acquire syntax are among the central questions in linguistics and cognitive science (Christiansen and Kirby, 2003; Nowak and Komarova, 2001; Fitch, 2000). Due to lack of direct evidence for language evolution, many studies employ evolutionary modeling (Christiansen and Dale, 2003; O’Grady and Smith, 2018) in order to gain insights into the putatively most difficult problem in science. While many studies achieved substantial progress (Nowak and Krakauer, 1999; Kirby, 2000; Griffiths and Kalish, 2007; Zuidema and de Boer, 2009; Perfors and Navarro, 2014; O’Grady and Smith, 2018), most models, however, are either theoretical or operate with highly simplified and abstract assumptions. In this paper, we model the first step in the evolution/acquisition of syntax—concatenation—with unsupervised deep neural networks that are among the most realistic models of human language learning and learn linguistic representations from raw spoken language inputs (Beguš, 2021a).
[…]
Concatenation in itself is not yet syntax. Many other systems, such as bird song (Berwick et al., 2011; Okanoya, 2013; Youngblood, 2024) or music (Patel, 2007) contain concatenation with no referential meaning. The degree to which these systems are analyzed in terms of syntactic processing is still disputed. In our proposal, meaning is modeled in a highly abstract way with binary codes. This is only an imperfect, but not insurmountable approximation of the actual referential meaning. Concatenating raw phonetic items into multiple-word outputs is thus the very first and most basic precursor to syntax, and not syntax itself. However, without this step, no further syntactic evolution could have been possible. Human syntax is also fully compositional. We show that our model not only spontaneously concatenates, but also does so in a compositional manner for at least a subset of lexical items, which presents an additional step towards the evolution of human syntax.
As our model of learning we adopt ciwGAN and fiwGAN (Beguš, 2021a) which model language as informative imitation, feature several advantages over competing computational model, and represent one of the closest approximations of human language learning using deep neural networks. The sounds of human speech are a measurable, physical property, but they also encode abstract linguistic information such as syntactic, phonological, morphological, and semantic properties. It has been shown that the model learns symbolic-like rule-like linguistic representations from raw speech (Beguš, 2021a, b, 2020, 2022) which illustrates that discrete mental representational units can emerge from continuous physical properties such as sound in deep neural networks. The model features several aspects of human spoken language learning that other models lack: learning by imitation, learning through the production-perception loop, and communicative intent.
Prior models of human syntax are predominantly text-based and mostly lack explicit mechanisms that can be directly paralleled to language processing in the human brain. Hearing human learners, however, acquire syntax from acoustic inputs and process it with biological neural mechanisms. Few studies model syntax from speech with artificial neural networks. Previous work has shown that speech-trained models can feature syntactic representations, but the models featured specific syntactic mechanisms (Lai et al., 2023) or were pre-trained (Singla et al., 2022). These architectures do not model the emergence of syntax.
Here, we propose that we can model a shift from single unit system to a concatenative and partially compositional system using ciwGAN. We model how compound signals or concatenated words can arise spontaneously in deep neural networks trained on raw speech in a fully unsupervised manner. We thus also provide evidence suggesting that domain-general learning models (such as CNNs) spontaneously concatenate signals in a learning setting that involves informative imitation: learning by imitation and with the production-perception loop from raw spoken language data (ciwGAN). The paper thus introduces a model of the first step in the development from holistic non-syntactic signals to compositionally concatenated signals that represents the precursor to syntactic processing. We also show that the observed spontaneous concatenation is not an idiosyncratic property of a single model, but emerges in several replicated models as well as in models that features several changes in the architecture.
In Section 4.1, we provide a plausible neural explanation for the concatenative behavior called the disinhibition which suggests that such modeling can be useful for generating and simulating predictions for neural processing of speech. We also show that the networks exhibit traces of compositionality (Section 3.6), which represent an additional step toward the evolution of human syntax. In Section 4.2, we formalize the interaction between latent variables and connect it to existing mathematical treatments of the Merge operation.
2 Methods
To test whether CNNs can spontaneously concatenate lexical items, we conduct two sets of experiments. In the first set of experiments (four one-word experiments), we train the networks on single-word inputs and test the networks for concatenated outputs. In the second experiment, we train the networks on one-word and two-word inputs (the two-word experiment) and withhold a subset of two-word combinations. We then test whether words can be embedded into novel unobserved combinations in the output. Such a design also mimics one-word and two-word stages in language acquisition (Berk and Lillo-Martin, 2012).
Because the networks are trained in the GAN setting, the Generator never accesses the training data directly, but generates innovative outputs. It has been shown that GANs innovate in highly interpretable ways that produce novel words or sound sequences (Beguš, 2021a). Here we test whether innovations can produce spontaneously concatenated words.
2.1 The model
We train the ciwGAN and modified fiwGAN models (Beguš, 2021a). ciwGAN/fiwGAN models are information-theoretic extensions of GANs (based on InfoGAN, WaveGAN and DCGAN; Chen et al. 2016; Donahue et al. 2019; Radford et al. 2015) designed to learn from audio inputs.
The ciwGAN/fiwGAN architectures involve three networks (Figure 1 and Table 3): the Generator that takes latent codes $c$ (either one-hot vectors or binary codes) and random latent space variables $z$ ($z\sim\mathcal{U}(-1,1)$) and through five or six upconvolutional layers generates 1.024s or 2.048s audio with 16 kHz sampling rate (16,384 or 32,768 samples). The audio is then fed to the Discriminator, which evaluates realness of the output via the Wasserstein loss (Arjovsky et al., 2017). The unique aspect of the ciwGAN/fiwGAN architecture is a separate Q-network, which is trained to estimate the Generator’s hidden code $c$. In the ciwGAN architecture, the latent code $c$ is a one-hot vector; in the fiwGAN architecture the latent code is a binary code. It has been shown that even if there is a mismatch between the number of lexical items and the number of classes, lexical learning nevertheless occurs in the models (Beguš, 2021a).
Figure 1: The architecture of ciwGAN used in the two-second one-word experiment.
During training the Generator learns to generate data such that it increases the Discriminator’s error rate and decreases the Q-network’s error rate. In other words, the Generator needs to learn to encode unique information into its acoustic inputs, such that the Q-network is able to decode unique information from its generated sounds. The training between the Generator and the Q-network mimics the production-perception loop in speech communication: after training, the Generator learns to generate individual words given a latent code $c$ (Beguš, 2021a) and the Q-network learns to classify unobserved words with the same corresponding codes (Beguš and Zhou, 2022c). Since learning is completely unsupervised, the Generator could in principle encode any information about speech into its latent space, but the requirement to be maximally informative causes it to encode linguistically meaningful properties (both lexical and sublexical information; Beguš and Zhou 2022c). Such a setting not only replicates the production-perception loop, but is also one of the few architectures featuring traces of communicative intent (between the Generator and the Q-network). Unlike in generative models trained on next sequence prediction or data replication where no communicative intent exists, the training objective between the Generator and the Q-network is to increase mutual information between the latent space and the data such that the Q-network can retrieve the information (latent code) encoded into the speech signal by the Generator.
CiwGAN and fiwGAN have been shown to be highly innovative in linguistically interpretable ways (Beguš, 2021a, b). For example, the Generator produces new words or new sound sequences that it never accesses during training. Crucially, the Generator never directly accesses the data: it learns by generating data from noise such that the Discriminator fails to distinguish real and generated data. In this respect, it mimics learning by imitation in human language (rather than replication as is the case with variational autoencoders).
Three models are trained for the purpose of this paper, and two models are pretrained from Beguš (2021a). The models were trained for approximately 16 hours on a single GPU (NVIDIA 1080ti). For guidelines on the choice of the number of steps during training, see Beguš (2020). We use the TIMIT (Garofolo et al., 1993) database for training. The number of parameters is given in A. We take the standard hyperparameters (from Donahue et al. 2019 and Beguš 2021a). Because the outputs are salient, we combine transcriptions performed by the authors with transcriptions performed by Whisper (Radford et al., 2022). Generated audio files and models’ checkpoints are available at this open source repository: https://osf.io/przuq/.
2.2 Data
The training dataset consists of sliced lexical items from the TIMIT database of spoken English (Garofolo et al., 1993), such that each item is a single spoken word.
In order to test robustness of spontaneous concatenation across different models and settings, we conduct four experiments with varying parameter choices. We alter the number of training steps, training words, word-types, and tokens, randomness of padding, length of output and model type. The experiments are summarized in Table 1.
In Experiment 1 and 2, the one-second one-word experiment, we test two pre-trained models from Beguš (2021a) for signs of concatenation (in Experiment 2, we use a version of the pre-trained model with fewer training steps). As specified in Beguš (2021a), the 5-layer Generator outputs only 1.024s of audio and data is never left-padded (only right padded) which controls for the effect of padding on concatenation. It is possible that the Generator would learn that silences can occur on both sides of the lexical item in training, and then inverse silences with lexical items which would result in concatenation. The right-padding experiments control for this possibility.
The model in Experiment 1 uses 5 lexical items with  600 tokens each: oily, rag, suit, year and water. The 10-word model (Experiment 2) is trained on dark, water, oily, year, greasy, like, carry, ask, rag, suit. The smaller number of lexical items chosen for training is advantageous from the perspective of modeling language acquisition as it mimics the stage when language-acquiring children have fewer lexical items in their inventory.
In Experiment 3 (simple two-second one-word), we replicate the results with another one-word experiment trained on box, greasy, suit, under, and water. Here, each item is represented by a single token (to test how the model performs with less variability during training) and randomly padded with silence to a length of 2s to produce 100 distinct data points for each class, for a total of 500 data points used in training (the two-second one-word experiment).
We perform Experiment 4 (complex two-second one-word) by training a model where each lexical item is represented by approximately 600 distinct tokens, as found in the TIMIT dataset, to test the ability of the model to learn in a more complex environment. Again, for this experiment each data sample is randomly padded with silence to a length of 2s.
Experiment Words per file Length of output Padding Types Tokens Steps Model
1 1 1.024s right 5 $\sim$600 8,011 ciwGAN
2 1 1.024s right 10 $\sim$600 7,678 ciwGAN
3 1 2.048s random 5 1 8,956 ciwGAN
4 1 2.048s random 5 $\sim$600 18,842 ciwGAN
5 2 2.048s random 3 1 9,166 fiwGAN
Table 1: An overview of parameters used in the four experiments.
In Experiment 5 (two-second two-word), we use 3 lexical items with a single token each: greasy, suit, and water. 100 data points each 2s in length are generated in an analogous process to the first experiment, but for each combination of two items (i.e. greasy, suit, and water alone, greasy followed by water, water followed by greasy etc.). However, we withhold the combination suit/greasy, such that they do not appear together in the training set in any order, to produce a final training set of 700 data points.
For the one-word experiments, we use the ciwGAN model with one-hot encoding and five levels, such that each of the five unique words can be represented with a unique one-hot vector. In the two-word experiment, we use a modified fiwGAN (binary code). The binary code is limited to three bits, but each code can have up to two values of 1 (e.g. [1,0,0] and [1,1,0]).
We use Experiment 3, 4, and 5 to show that concatenation happens overwhelmingly in the negative values. The advantage of these models is that they have random padding, which results in a sufficient audio quality of concatenated outputs for automatic Whisper evaluations. The disadvantage of these models is that the networks can use the latent code as positional encoding of outputs rather than lexical learning. We use Experiments 1 and 2 to show that lexical learning happens in the code space $c$ and that the same code space results in concatenation in negative values (parallel to Experiments 3, 4, and 5). Additionally, Experiments 1 and 2 show traces of compositionality in concatenation of lexical items based on the code variables $c$ which strengthens our argument that these models are mimicing initial stages of the evolution of syntax. In these experiments, we show that the models learn to represent words in a fully unsupervised way with latent codes and, at the same time, learn to concatenate them compositionally in a subset of word pairs. Based on internal intepretability of the models (as proposed by Beguš 2020, 2021a; Beguš and Zhou 2022b, a), we outline neural mechanisms behind the compositional concatenation and formalize the development towards compositional syntax based on a plausible neural pathway called disinhibition.
3 Results
To test whether the models can spontaneously concatenate, the networks are trained for 8,011, 7,678 (pretrained one-second one-word models), 8,956 (simple two-second one-word), 18,842 (complex two-second one-word), and 9,166 (two-word fiwGAN) steps and analyze generated data. We use the technique proposed in Beguš (2020) to analyze the relationship between linguistically meaningful units and latent space. According to this technique, setting individual latent space variables to values outside of the training range reveals the underlying linguistic value of each variable.
3.1 One-word model
In the one-second one-word model (Experiment 1 in Table 1), the Generator learns to associate each unique one-hot code with a unique lexical item in a fully unsupervised and unlabeled manner (Beguš, 2021a). The Generator’s input during training is a one-hot vector with values 0 or 1. For example, the network learns to represent suit with $[1,0,0,0,0]$. To test this observation, we set the one-hot vector to values outside the training range (e.g. $[5,0,0,0,0]$), which overrides lower-level interactions in the latent space (Beguš, 2020). This causes the Generator to output suit at near categorical levels (9 times out of 10), revealing the underlying value of the code. This further reveals that $[0,1,0,0,0]$ encodes year (8 times out of 10) and $[0,0,1,0,0]$ encodes water (10 times out of 10).
In addition to lexical learning, we observe a robust pattern in the pretrained one-word one-second model: the networks trained on one-word inputs generate two-word outputs when the one-hot values are set to negative values outside of the training range. For example, when the latent code is set to $[0,-2,-2,-2,0]$, the Generator consistently outputs a two-word output suit year (8 times out of 10). For $[-3,-2,-3,-2,2]$, the network consistently outputs rag year (8 times out of 10; Figure 2). These concatenations occur despite the fact that the training data is always left-aligned, the Generator never accesses the data directly and the Discriminator only sees single words.
Figure 2: The suit year (left) output and the rag year (right) from the one-second one-word model. All spectrograms are created in Praat (Boersma and Weenink, 2015).
To show that this concatenation is not an idiosyncratic property of one model and that it is indeed the negative values that encode concatenated outputs, we also analyze the simple two-second one-word model (Experiment 3 in Table 1). The inputs to the Discriminator in this case are also single words only, but they are longer (2s) and randomly padded with silence on the left and right. While high positive values occasionally yield two-word outputs in this model, negative values are consistently associated with two-word outputs. For example, $[-50,-50,0,-50,0]$ (with extreme values) consistently encodes box greasy (9 times out of 10), and $[-50,-50,-50,0,0]$ consistently encodes greasy under (10 times out of 10). Positive values of the same codes with extreme values (50) produce completely unintelligible outputs (noise).
The same outcomes are observed in the Experiment 4 (in Table 1) where the models are trained on approximately 600 token for each lexical item (for a quantitative analysis of this experiment, see Section 3.2 and Figure 4).
In addition to several two-word concatenated outputs, the network even occasionally generates a three-word concatenated output box under water (Figure 3) for the latent code $c$ with all negative values $[-3,-1,-1,-1,-1]$ (2 times out of 10).
The disadvantage of the randomly padded models is that the one-hot code $c$ encodes the position of the words in the output rather than words itself, which means that for the purpose of testing compositionality with lexical learning and concatenation, we focus on the right-padded model only, where robust lexical learning in the one-hot codes $c$ has been documented (Beguš, 2021a).
Figure 3: The three-word concatentated output box under water. Independently, the second word (under) is somewhat difficult to analyze, but given only five training words, it is clearly the closest output to under.
3.2 Statisical analysis
Figure 4: The average number of words generated by the two one-word two-second models are plotted as a function of the first 3 bits of the latent code. The remaining two bits of the latent code are maintained at a value of 0. As in the other trials, each bitstring was tested with 10 sets of latent space values.
Raw data in Figure 4 shows that the simple one-word two-second model (Experiment 3) generates between 0 and 3 words for each $z$ vector, generating substantially more words at negative bit values. To test the observation that negative values correspond to concatenated outputs, we systematically generate samples where each of the 5 bits is interpolated from the set of values $\{-6,-3,0,3,6\},$ and all permutations of bits are tested ($5^{5}$). Each permutation is tested with 10 sets of latent space values (constant across different permutations), for a total of $31,250$ samples. We automatically annotate our outputs using a pretrained Whisper-small model (Radford et al., 2022), which we fine-tune using a dataset consisting of 400 model outputs manually annotated by the authors. Whisper is used to annotate for single (failure; 0) vs. multiple (success; 1) word outputs. We estimate Whisper’s accuracy on labeling single- and multiple-word outputs at 88% (based on author’s annotations; Table 5). The data were fit to a logistic regression mixed effects model (details of the model in Figure 5). There is a clear and significant negative relationship between the sum value of bits and the probability of a multiple-word output ($\beta=-0.09,z=-54.42,p<0.0001$; Figure 5). The model with interactions between individual bits also shows that negative values correspond to significantly higher proportion of two-word outputs (Figure 5, Table 4).
We perform a similar analysis on the complex one-word two-second model (Experiment 4 in Table 1), where we systematically generate samples where each of the 5 bits is interpolated from the set of values $\{-4,-2,-1,0,1,2,4\}$, and all permutations of bits are tested $(7^{5}),$ Again, each permutation is tested with 10 sets of latent space values (constant across different permutations), for a total of $168,070$ samples. All samples were subsequently transcribed using the same fine-tuned Whisper model used in the previous experiments, and the results suggest a similar correlation between negative code values and the proportion of two-word outputs (Figure 4).
The data of this experiment were fit to a logistic regression model which again shows a significant relationship between the proportion of multiple-word outputs (the dependent variable) and the sum of all the latent code bits (the independent variable) ($\beta=-0.14,z=-167.5,p<0.0001$). All details of the model and plotted estimates are in Figure 13.
Figure 5: (top) Predicted values of the logistic regression mixed effects model with the proportion of two-word outputs (one-word output = failure, two-word output or more = a success) as the dependent variable and sum of bits 1–5 as the predictor with 95% confidence limits (Experiment 2 in Table 1). Outputs with no transcribed words were removed. The random effect structure involves random intercepts for each of the 10 unique random latent space samples $z$ as well as random slope for sum of bits. The values were obtained with the effects package (Fox and Weisberg, 2019; Fox, 2003). (bottom) Predicted values for one (bit1 * bit2 * bit5) of the many interactions of the logistic regression mixed effects model with the proportion of two-word outputs (one-word output = failure, two-word output = a success) as the dependent variable and individual bits with all interactions (including the five-way interaction) as the predictors with 95% confidence limits. The random effect structure involves only the random intercept for each of the 10 unique random latent space samples $z$. The values were obtained with the effects package (Fox and Weisberg, 2019; Fox, 2003). Model estimates are given in Table 4. Not all interactions show the same effect.
The concatenated words are also well distributed across word types. We utilize the same dataset of $168,070$ outputs (with all permutations of code values $\{-4,-2,-1,0,1,2,4\}$) transcribed of by the fine-tuned Whisper as was used for the regression model to estimate the distribution of words across the positive and negative latent code values. Fine-tuned whisper has approximately 54% match rate with our human evaluations, but this criterion is conservative as only perfect matches are count as success. In single word outputs, the word water is overrepresented with the rest of the lexical inventory relatively equally represented. In concatenated word outputs, the distribution of word types is also relatively uniform. The five training words are well represented in both the first and second positions, which suggests that word concatenation is not an idiosyncratic property of a few word types or an acoustic artefact. Figure 6 illustrates the most common types for single-words and for concatenated outputs.
Figure 6: Counts of word types as transcribed by fine-tuned Whisper. (a) Counts of 7 most frequent outputs that were transcribed as including a single word by the fine-tuned Whisper. (b) Counts of 20 most frequent outputs that were transcribed as including two or more words by the fine-tuned Whisper.
3.3 Two-word model
In the two-word experiment, the models get one-word and two-word inputs (thus mimicking the two-word stage in language acquisition). The models are only trained on three words and their combinations, except for the suit/greasy combination withheld.
In the fiwGAN two-word model (Experiment 5 in Table 1), the Generator consistently outputs the unobserved greasy suit for $[15,0,0]$ (17 times out of 20), which suggests the network learned this unobserved combination as one of the possible sequences and encoded it with a one-hot value. To evaluate the presence of the withheld greasy suit pair systematically, we conduct an exploration of the latent space (all permutations where each variable takes on values in $\{-13,-8,-3,-2,-1,0,1,2,3,8,13\}$) in Figure 7. The withheld suit/greasy combinations occur most often for codes around [-10,0,0], or otherwise when at least one bit is negative (Figure 7), suggesting that the model has associated the unobserved suit/greasy combination with that code. It appears that the negative values of the latent code $c$ again encode unobserved novel combinations.
Figure 7: The frequency of observed suit/greasy pairs is plotted as a function of the latent code, for the fiwGAN two-word model. Samples were generated by varying each bit between the values [-13, -8, -3, -2, -1, 0, 1, 2, 3, 8, 13], taking all possible permutations ($11^{3}$). Each bitstring was tested with 10 sets of latent space values, and annotations were performed automatically using Whisper.
3.4 Repetition
In addition to two-word concatenation and embedding of words into novel combinations, we also observe outputs with repeated words in all our trained models. The training data never includes the same word repeated, yet the models frequently include repeated words. For example, the two-second one-word model consistently outputs greasy greasy for $[0,0,-40,0,0]$ (7 times out of 10; Fig. 8). Of the 31,250 samples annotated in the analysis, 2753 (8.8%) contained repeated words, suggesting that the model systematically learned this behavior. This is significant because repetition or reduplication is one of the most common processes in human language and language acquisition (Berent et al., 2016; Dolatian and Heinz, 2020).
Figure 8: The suit greasy water output from the two-second two-word model (top). The greasy greasy output from the two-second one-word model (bottom).
3.5 Single batch model
We consider the possibility that spontaneous concatenation arises solely due to the models being trained in a batched training configuration, where different samples in the same batch could be concatenated when their gradients are combined into a single training update. To demonstrate that spontaneous concatenation occurs even in the absence of batching, we train an additional one-word model using the same dataset and model parameters as the complex two-second one-word model, but with the batch size set to $1$ We trained this model for 41,922 steps. It is well-known that batch size can affect training outcomes (Lin, 2022). Because of the small batch size, the words generated by this model are qualitatively poorer which prevents an automated Whisper analysis of the outputs. However, we continue to consistently observe multiple-word outputs in the negative latent space, indicating that this behavior is not solely attributable to batching. We randomly generated 10 examples from this model with the latent code set at [-1, -1, -1, -1, -1]. All 10 examples feature two or more word sequences, although the audio quality is too low to identify the words. Waveforms and spectrograms of the ten examples are given in Figure 14.
3.6 Towards compositionality
Compositionality has long been focus of computational models (Kirby 2001; Andreas 2019; Chaabouni et al. 2020; Rita et al. 2022; for a review, see Rita et al. 2024). Most evaluated models, however, are trained on already compositional data or operate with symbolic units rather than with raw data. Our model is trained on non-compositional single unit words. Concatenation and compositionality need to self-emerge in our setting.
Experiments in Section 3.1 suggest that negative values in the latent space result in spontaneously concatenated outputs. Here, we show that these spontaneously concatenated outputs show traces of compositionality.
The ciwGAN models have been shown to learn lexical items (Beguš, 2021a). In a fully unsupervised manner, the network learns to associate unique codes with unique lexical items. For example, when a ciwGAN network is trained on approximately 600 tokens of five lexical items, it learns to associate [1, 0, 0, 0, 0] with the word suit, [0, 1, 0, 0, 0] with year, [0, 0, 1, 0, 0] with water, [0, 0, 0, 1, 0] with oily, and [0, 0, 0, 0, 1] with rag (Beguš, 2021a). Lexical learning thus emerges in a fully unsupervised way, only from the requirement that the networks can recover the latent code from synthesized audio.
Single lexical items are represented with unique one-hot codes can be causally forced in the output at near categorical levels. For example, setting the code to [0, 0, 0, 0, 2], the network outputs rag in 98% of outputs.
Exploration of the negative extreme values in the one-hot code $c$ shows that some values represent a specific single word or a word concatenated with another, non-specific word that is difficult to acoustically identify. For example, in Experiment 1, the model consistently outputs the word year together with another non-specific word when its latent code is set to [0, -10, -10, 0, 0]. Figure 9 illustrates one such example. We generate five other outputs (audio samples available at the data repository link), but they are nearly identical to the one displayed in Figure 9.
Figure 9: Waveform and spectrograms (0–5kHz) of three generated outputs with latent codes [0, -10, -10, 0, 0], [0, -10, -10, 0, 5], and [5, -10, -10, 0, 0].
The first word in this output (Figure 9) can be transcribed as rag but its quality is low. The second word can be transcribed as year. It appears that [0, -10, -10, 0, 0] represents the word year concatenated with another word with a weaker audio profile.
This becomes apparent when we combine [0, -10, -10, 0, 0] with the code that represent the word suit ([1, 0, 0, 0, 0]) and set both to extreme values ([5, -10, -10, 0, 0]). When the latent code $c$ is compositionally combined in such a way, we get a clear case of a compositional concatenation in the output: suit year. We generate 5 outputs and they are all near identical and contain the suit year concatenation. When we combine [0, -10, -10, 0, 0] with the value that represents rag, we get a clear output containing rag year. We can thus causally intervene and compositionally force the desired lexical concatenated items in the outputs. This happens in the model that was only trained on single right-padded word inputs. While such a clear causal relationship between the compositionality of the latent code and the compositionality of the concatenated outputs exist for a subset of codes and not all codes, this is clearly a step towards a compositional representation of concatenation.
Compositionality is even more pronounced in the second right-padded one-word model (Experiment 2). The model, trained on ten words, learns to represent words with unique one hot codes. For example, [1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0] represents the word dark, etc. Again, setting the values of the code to extreme values (in line with Beguš 2020, 2021a) reveals the underlying value of each code and results in a single lexical item in the output at near categorical levels.
We replicate the finding from Experiment 1: setting the latent code to extreme negative values in a subset of cases results in a specific word either surfacing alone or concatenated with another word with a low acoustic profile. For example, in about half of the outputs, the code [0, 0, -1.5, 0, 0, 0, 0, 0, 0, 0] results in a single word greasy with a relatively low, but identifiable quality (in about one half of the outputs are not identifiable because of poor acoustics). When this code is compositionally combined with extreme positive values of codes that represent other words, we get predictable concatenated words of greasy together with the words that a specific code represents. This strongly suggests [0, 0, -1.5, 0, 0, 0, 0, 0, 0, 0] represents greasy and that compositionally combining this code with codes that represent other words lead to a predictable concatenation.
For example, code [0, 0, 0, 1, 0, 0, 0, 0, 0, 0] represents year. A combination of this word plus the negative code representing greasy gives a clearly identifiable year greasy ([0, 0, -6.5, 4, 0, 0, 0, 0, 0, 0]) in 10/10 generated outputs. A combination with the codes representing ask, suit, and water results in ask greasy, suit greasy, and water greasy in 10/10 or 9/10 outputs. Table 2 illustrates the full compositionality of the concatenated outputs and Figure 10 illustrates the causally compositional outputs with waveforms and spectrograms.
Negative code Word Positive Code Word Positive & Negative Combined Concatenated Words
[0, 0, 0, 1, 0, 0, 0, 0, 0, 0] year [0, 0, -1.5, 0, 0, 0, 0, 0, 0, 0] greasy [0, 0, -6.5, 4, 0, 0, 0, 0, 0, 0] year greasy
[0, 0, 0, 0, 0, 0, 0, 0, 0, 1] suit [0, 0, -1.5, 0, 0, 0, 0, 0, 0, 0] greasy [0, 0, -5.0, 0, 0, 0, 0, 0, 0, 3] suit greasy
[0, 0, 0, 0, 0, 0, 0, 1, 0, 0] ask [0, 0, -1.5, 0, 0, 0, 0, 0, 0, 0] greasy [0, 0, -8.0, 0, 0, 0, 0, 3, 0, 0] ask greasy
[0, 1, 0, 0, 0, 0, 0, 0, 0, 0] water [0, 0, -1.5, 0, 0, 0, 0, 0, 0, 0] greasy [0, 1, -8.0, 0, 0, 0, 0, 0, 0, 0] water greasy
[0, 0, 0, 0, 0, 0, 0, 0, 0, 1] suit [0, 0, 0, 0, 0, 0, -8.0, 0, 0, 0] X carry [0, 0, 0, 0, 0, 0, -12, 0, 0, 7] suit carry
[0, 0, 0, 0, 0, 0, 0, 1, 0, 0] ask [0, 0, 0, 0, 0, 0, -8.0, 0, 0, 0] X carry [0, 0, 0, 0, 0, 0, -24, 5, 0 ,0] ask carry
Table 2: Compositionality of codes and their outputs.
Figure 10: Waveforms and spectrograms (0-5 kHz) of five generated outputs with latent code variables as described in Table 2. The first output (from left) represent greasy with negative values of the latent code $c$. The following outputs have compositional values of the latent code $c$: the negative value associated with greasy is combined with the positive value that represent individual lexical items. The outputs contain compositionally concatenated words with predictable word-pairs based on the latent code.
Compositional outputs, while not available for all lexical items, are not limited to greasy only. For example, code [0, 0, 0, 0, 0, 0, -8, 0, 0] represents carry in concatenation with a word that has unrecognizable acoustics (potentially close to ask). When this code is causally combined with the code representing suit, the output is a concatenated suit carry. While the acoustics are not as clear as in the greasy examples, the concatenations are clearly compositional. When the code representing concatenated carry is compositionally combined with the code representing ask, the output results in a clearer ask carry. Table 2 illustrates this compositionality.
All generated outputs with the exception of the one representing greasy are nearly identical across 10 generated repetitions and are available in the data repository. Spectrograms in Figures 10 and 15 are thus representative of the outputs’ acoustic content.
Both independently trained models in Experiment 1 and 2 thus learn lexical items and learn to concatenate them compositionally at least in a subset of possible combinations.
3.7 Limitations
This paper models concatenation of acoustic lexical items with traces of compositionality. Syntax is substantially more complex than concatenation and compositionality (Berwick and Chomsky, 2019). Meaning is in our models represented abstractly with a one-hot encoding. While one-hot code can abstractly represent multimodal meaning, a more realistic model would have to contain referential meaning.
We also train the network on a relatively small number of lexical items (5) and a small number of tokens (from one to approximately 600). The small number of lexical items is representative of the earliest stages of language acquisition when the number of lexical items is highly limited (Bates et al., 1994). Human syntax, however, operates on a substantially larger number of elements. How this approach scales up is left for future work.
4 Discussion
4.1 Artificial disinhibition
Here we propose a potential neural explanation for why negative values result in outputs with concatenated words. We propose that spontaneous concatenation results from artificial disinhibition. Disinhibition (or excitation that results from inhibition of inhibitory neurons; Pi et al. 2013; Pfeffer et al. 2013) is a well-established biological neural process. Negative values in the latent space function as inhibitory neurons. If inhibitory neurons operate on another set of inhibitory neurons, the result is excitatory (the so-called disinhibition; Pi et al. 2013; Pfeffer et al. 2013).
The first layer in our convolutional Generator is the fully connected layer. In this layers, all 100 latent variables (code variables $c$ and random variables $z$) are connected to all variables in the fully connected layer with weights that get updated during training. The fully connected layer is of the shape $1024\times 16$, which means that there are 1024 time series of length 16. The 16 dimensions represent the time domain and get transformed to the time dimension in the final output via the convolutional layers (Figure 11).
The fully connected layer needs to include both excitatory artificial neurons (or positive weights that generate words) as well as inhibitory artificial neurons (or negative weights that generate non-activity or silence). The non-activity or silences are likely generated by negative weights which results in negative values of the fully connected layer. When negative values in the latent space get multiplied with negative weights, however, the resulting output in the fully connected layer will involve positive values. This inhibition of inhibitory weights (or disinhibition) is the mechanism whereby the Generator can generate multiple words in the output. This process mimics the biological disinhibition in human neuroscience.
Figure 11: The structure of the Generator. Each variable in the latent space (5 code variables $c$ and 95 uniformly distributed variables $z$) are fully connected to all dimentions of the fully connected layer ($16\times 1024$). The figure only illustrates a few connections. The fully connected layer contains 1024 time series data with the dimension of 16. The 16 dimensions correspond to the time domain in the audio output. The output of the fully connected layer gets upsampled to audio output in the final layer via 5 convolutional layers. The six time series data above the Fully connected layer box illustrate the first five out of 1024 time series data. In this paper, we analyze averages of 1024 time series data in the Fully connected layer.
That inhibitory neurons play an important role in human language processing in the brain has been well established. The role of inhibitory neuron has primarily been established for segmentation (Giraud and Poeppel, 2012). Segmentation is closely related to concatenation, which lends some biological plausibility to our proposal.
Figure 12 illustrates the pathway to how negative values (or disinhibition) result in concatenated outputs. We plot ten outputs from the fully connected layer for each of the five one-hot codes with positive values (with the value of the code set to 10, e.g. [10, 0, 0, 0, 0]) and the corresponding negative values (e.g. [-10, 0, 0, 0, 0]) of the model used in Experiment 1. The fully connected (Dense) layer consists of 1024 instances of 16-valued outputs. The 16 values represent the time time domain and get transformed to the actual temporal dimension in the generated audio through the convolutional layer. In line with Beguš and Zhou (2022b, a), we average each output in the time domain, which yields a single 16-dimension output that summarizes neural activity in the deepest layers in the convolutional network. We plot both pre-ReLU and post-ReLU values.
That words are represented by unique shapes in earlier convolutional layers has been shown previously by Beguš and Zhou (2022a). Negative values in the latent space cause the fully connected layer to have inverse of the positive values (or disinhibitory values) as is clear from Figure 12. After the ReLU activation (which itself is a biologically plausible operation), the negative or disinhibited values (inverse of the positive values) can cause multiple shapes rather than a single shape with positive values. The shapes in the deeper layers get then transformed into distinct lexical items through a series of convolutional layers (Beguš and Zhou, 2022b, a).
Figure 12: Averaged values of the Dense layer output before (Avg. Dense) and after ReLU (Avg. ReLU) for positive and negative extreme values of the latent code $c$. Each latent code (e.g. [10, 0, 0, 0, 0] or [-10, 0, 0, 0, 0]) was sampled randomly ten times.
This line of reasoning thus uncovers a possible neural mechanism behind the precursors of syntactic concatenation or even the simplified Merge operation (Chomsky, 2020). In the initial stage, the concatenated output can be represented as a disinhibition, which generates excitatory activity of two or more words as part of a single output. In other words, a concatenated output has a unique representation: disinhibition that results in activation of two words. This point represents a non-compositional concatenative stage in the evolution/acquisition of syntax.
As suggested by our experiments in Section 3.6, the next development towards compositionality occurs when the disinhibited values start representing a single word concatenated with another word. At this point, a combination of excitation that activates one word and disinhibition that activates another word can result in a causally predictable output that compositionally concatenates the two words.
Disinhibition outlines a a potential artificial and biological neural pathway towards concatenated outputs and compositionality in human syntax. Testing this hypothesis from neural and evolutionary standpoint is left for future work.
4.2 Formalizing disinhibition
Here, we outline how a precursor to the Merge operation can empirically and numerically emerge in a fully connectionist model. Chomsky (2020) and Marcolli et al. (2023), for example, define Merge ($\mathfrak{M}$) as in 1, where $\alpha$ and $\beta$ are arguments.
$\mathfrak{M}(\alpha,\beta):=\{\alpha,\beta\}$ (1)
Under the minimalist syntactic approach, the $\{\alpha,\beta\}$ is compositional. In our case, the apparent $\{\alpha,\beta\}$ is represented idiosyncratically as the disinhibitory neuron. The disinhibitory neuron causes the concantenated shapes of two or more words, which are executed by the Generator as two words in the output. We simplify $\alpha$ and $\beta$ to represent lexical items with some associated meaning in the form of one-hot codes. According to the formalism, we can represent the Disinhibition ($\mathfrak{D}$) as:
$\mathfrak{D}(\gamma)=\{\alpha,\beta\}$ (2)
This step does not represent the full extent of compositional concatenation, because there is no relationship between $\gamma$ and the merged $\{\alpha,\beta\}$. However, this step represents the first operation required in the development from the non-concatenative simple signals to complex compound signals of human syntax.
Based on our experiment in Section 3.6, it appears that the step from non-compositionality toward compositionality occurs through an intermediate stage when the negative, disinhibitory values start representing a specific lexical item concatenated with another non-specific word. We have shown that a subset of disinhibited values consistently represent a single lexical item or the specific lexical item concatenated with another non-specific item. Such disinhibitory values are formalized in 3.
$\mathfrak{D}(\gamma)=\{\alpha,\}$ (3)
Compositional concatenation can now be achieved by joining the values of $\mathfrak{D}(\gamma)$ and $\beta$.
$\mathfrak{C}(\mathfrak{D}(\gamma),\beta)=\{\alpha,\beta\}$ (4)
In our model, $\mathfrak{C}$ is a simple operation achieved by simultaneously setting two or more values of the latent code $c$ (that represents the meaning of the words) to extreme values as outlined in Beguš (2020, 2021a). Because the relationship between $\mathfrak{D}(\gamma)$ and $\beta$ exists, 4 is compositional.
While structurally human syntax is fully compositional, we have little evidence that neural processing of language is compositionaly represented at the single neuron level in the human brain. The proposed model thus also allows simulating single-neuron effects on compositional syntax and generating predictions that can be tested in the human brain. The advantage of this model is that it operates with a fully continuous data, which is a more realistic approximation of human syntactic processing—human syntax is processed from raw unsupervised spoken language. All other models operate exclusively at the symbolic level.
4.3 The role of imitation
While disinhibition is a likely explanation for the concatenated outputs in the negative latent space, it needs to be noted that concatenated outputs happen also when the latent code values are positive (Figure 4).
The training data never provides evidence for concatenated outputs. It should be easy for the Discriminator to distinguish single- from multiple-word outputs and estimate multiple-word outputs as diverging from the real data distributions. One possible reason for why spontaneous concatenation is so common in GANs and why it occurs also in the positive values of the latent space is that unlike any other architecture, GANs learn primarily by imitation and they do not replicate data, but rather learn to construct data by innovation. In other words, in the GAN architecture, the agent that generates data (the Generator) never accesses training data directly, but learns to construct data from noise in the latent space by generating data such that another agent is unable to distinguish it from training data — effectively forcing the imitative principle. Imitation is a prominent facilitator of language acquisition in humans, especially in early spoken language learning (Clark, 1977; Kuhl and Meltzoff, 1996; Masur and Rodemaker, 1999; Rasilo and Räsänen, 2017).
It is possible that the imitative and innovative design of GANs mimics the necessary conditions for the switch from simple to compound signals in humans. Learning by imitation and innovation is a crucial condition in human language. It is possible that human language learners start uttering simple signals as concatenated and analyzing them as complex signals due to disinhibition as well as the imitative nature of learning. That syntactic processing develops at a relatively late stage and requires developments in the brain has been shown before (Hahne et al., 2004; Skeide et al., 2014). It has been shown elsewhere that the internal representations in ciwGAN/fiwGAN models closely match brain responses to sound of language in raw untransformed form (Beguš et al., 2023). These assumptions are, however, highly speculative at this point and require further evaluation.
5 Conclusion
Modeling the evolution from holistic single-call units into a compositional syntax is not a trivial task. How human syntax emerged or how humans evolved from the stage of holistic calls into a fully concatenative and then compositional syntax is an unsolved problem. While other studies offer proposals about how this shift may have happened (Jackendoff, 1999; Luuk and Luuk, 2014; Progovac, 2015), our paper shows that the shift spontaneously occurs in a computational model. Many studies exist that model evolution of syntax (Christiansen and Kirby, 2003), but the present study is, to our knowledge, the first that shows how concatenation and compositionality emerges spontaneously in a fully connectionist model that did not have any explicit predispositions for concatenation and that is trained on raw speech data.
Our results suggest that the Generator network in the ciwGAN architecture not only learns to encode information that corresponds to lexical items in its audio outputs, but also spontaneously concatenates those lexical items into novel unobserved two-word or three-word sequences.
We also provide evidence that the concatenation is at least partly compositional. The latent code is compositional. When negative values representing a word get combined with positive values representing another word, the output is a predictable concatenation of the two words. Our approach thus proposes that a causal compositionality for a subset of codes and words emerges in a fully connectionist model trained on raw speech.
We propose a neural mechanism for the emergence of compositional concatenation based on disinhibition. This explanation suggests that our modeling can be useful for generating testable biological and artificial neural prediction about language acquisition and evolution.
The ability of unsupervised deep neural networks trained on raw speech to compositionally concatenate words into novel unobserved combinations has far-reaching consequences. This means that we can model basic syntactic properties directly from raw acoustic inputs of spoken language, which opens up potential to model several other syntactic properties directly from speech with ciwGAN/fiwGAN.
From the perspective of evolution of syntax, the results suggest that a deep neural network architecture with no language-specific properties can spontaneously begin generating concatenated signals from simple signals. Precursors to compositionality emerge in a model that is trained on raw audio with no compositional training data points. The step from one-word stage to two-word stage is necessary both in evolution of human language as well as during language acquisition. Our second experiment mimics the two-word stage. We argue that unsupervised deep learning models not only concatenate single words into multi-word outputs, but are also able to embed words into novel unobserved combinations once the model is trained on multiple-word inputs.
Further research into the relationship between basic syntactic properties that spontaneously emerge in these fully unsupervised models trained on raw speech and the structure of the latent space has the potential to yield insights for the study of syntactic theory, language acquisition, language evolution, and neuroscience. By evaluating these models on syntactic properties of spoken language, we should also get a better understanding of computational limits of unsupervised CNNs trained on raw speech.
Finally, modeling syntax from raw speech with deep neural networks is informative not only for cognitive science and linguistics, but for machine learning research in general as well. Speech processing increasingly by-passes text (Lakhotia et al., 2021). Understanding syntactic capabilities of spoken language models can provide information for future architectural choices as natural language processing expands from text to raw audio modeling.
Acknowledgements
We would like to thank Bruno Ferenc Šegedin for his useful comments.
Author contributions
First and corresponding author: Gašper Beguš. G.B.: Conceptualization, Data curation, Formal analysis, Funding acquisition, Investigation, Methodology, Project administration, Resources, Software, Supervision, Validation, Visualization, Writing – original draft, Writing – review and editing. T.L.: Conceptualization, Data curation, Formal analysis, Investigation, Software, Visualization, Writing – original draft, Writing – review and editing. Z.W.: Conceptualization, Data curation, Methodology, Writing – review and editing.
Conflict of interest
The authors declare no conflicts of interest.
Appendix A Appendix
Generator Discriminator Q-Network
Layer Dimension Layer Dimension Layer Dimension
$z$ 100 $\times$ 1 input 32768 $\times$ 1 input 32768 $\times$ 1
fc + reshape 16 $\times$ 2048 conv0 8192 $\times$ 64 conv0 8192 $\times$ 64
upconv0 64 $\times$ 1024 conv1 2048 $\times$ 128 conv1 2048 $\times$ 128
upconv1 256 $\times$ 512 conv2 512 $\times$ 256 conv2 512 $\times$ 256
upconv2 1024 $\times$ 256 conv3 128 $\times$ 512 conv3 128 $\times$ 512
upconv3 4096 $\times$ 128 conv4 32 $\times$ 1024 conv4 32 $\times$ 1024
upconv4 16384 $\times$ 64 conv5 16 $\times$ 2048 conv5 16 $\times$ 2048
upconv5 32768 $\times$ 1 flatten + logit 1 $\times$ 1 flatten + logit 1 $\times$ 1
Table 3: The structure of the Generator, the Discriminator, and the Q-network in the two-second experiments (based on Donahue et al. 2019; Beguš 2021a).
Estimate Std. Error z value Pr($>$||||z||||)
(Intercept) -0.65 0.02 -36.7 0.0000
bit1 -0.21 0.00 -45.3 0.0000
bit2 -0.11 0.00 -24.4 0.0000
bit3 0.01 0.00 1.6 0.1013
bit4 -0.19 0.00 -41.2 0.0000
bit5 -0.17 0.00 -39.8 0.0000
bit1:bit2 0.03 0.00 30.5 0.0000
bit1:bit3 -0.02 0.00 -14.7 0.0000
bit2:bit3 -0.01 0.00 -7.0 0.0000
bit1:bit4 -0.00 0.00 -1.7 0.0911
bit2:bit4 0.03 0.00 28.2 0.0000
bit3:bit4 -0.02 0.00 -17.8 0.0000
bit1:bit5 0.02 0.00 22.7 0.0000
bit2:bit5 -0.03 0.00 -29.7 0.0000
bit3:bit5 -0.01 0.00 -7.2 0.0000
bit4:bit5 0.04 0.00 33.2 0.0000
bit1:bit2:bit3 -0.00 0.00 -13.9 0.0000
bit1:bit2:bit4 -0.01 0.00 -32.4 0.0000
bit1:bit3:bit4 0.01 0.00 39.1 0.0000
bit2:bit3:bit4 -0.00 0.00 -9.1 0.0000
bit1:bit2:bit5 -0.00 0.00 -1.9 0.0551
bit1:bit3:bit5 -0.00 0.00 -16.0 0.0000
bit2:bit3:bit5 0.00 0.00 0.5 0.5923
bit1:bit4:bit5 -0.00 0.00 -12.4 0.0000
bit2:bit4:bit5 -0.01 0.00 -19.8 0.0000
bit3:bit4:bit5 0.00 0.00 2.7 0.0060
bit1:bit2:bit3:bit4 0.00 0.00 2.5 0.0135
bit1:bit2:bit3:bit5 0.00 0.00 15.1 0.0000
bit1:bit2:bit4:bit5 0.00 0.00 17.8 0.0000
bit1:bit3:bit4:bit5 -0.00 0.00 -6.8 0.0000
bit2:bit3:bit4:bit5 0.00 0.00 3.5 0.0005
bit1:bit2:bit3:bit4:bit5 -0.00 0.00 -9.8 0.0000
Table 4: Estimates of the logistic regression mixed effects model with the proportion of two-word outputs (one-word output = failure, two-word output = a success) as the dependent variable and individual bits with all interactions (including the five-way interaction) as the predictors. The random effect structure involves only the random intercept for each of the 10 unique random latent space samples $z$.
Whisper annotated # of words
Human annotated # of words 1 2 3 Total
1 47 9 1 57
2 3 28 5 36
3 0 2 5 7
Total 50 39 11 100
Table 5: The accuracy of the Whisper speech-to-text model is assessed over 100 audio samples generated randomly using the two-second one-word ciwGAN model. Each sample was annotated manually by the authors and automatically using Whisper. For the classification of one word versus multiple words, Whisper agreed with the authors 88% of the time.
Figure 13: Predicted values of the logistic regression model with the proportion of two-word outputs (one-word output = failure, two-word output or more = a success) as the dependent variable and sum of bits 1–5 as the predictor with 95% confidence limits (Experiment 3 in Table 1). No random effects were included because mixed effects models result in singular fits. Outputs with no transcribed words were removed. The values were obtained with the effects package (Fox and Weisberg, 2019; Fox, 2003).
Figure 14: Ten randomly generated examples with the latent code [-1, -1, -1, -1, -1] of the single batch model.
Figure 15: Waveforms and spectrograms (0-5 kHz) of three generated outputs with latent code variables as described in Table 2. The first output (from left) represent carry with negative values of the latent code $c$. The word carry is preceded by another word which is difficult to identify, but has some acoustic properties of ask. The following outputs have compositional values of the latent code $c$: the negative value associated with carry is combined with the positive value that represent individual lexical items. The outputs contain compositionally concatenated words with predictable word-pairs based on the latent code.